OpenStudy (anonymous):

Fool's problem of the day, If \(a,b,c\) are in A.P and \(a^2,b^2,c^2\) are in G.P. If \( a

OpenStudy (lgbasallote):

a1 = a a2 = a + (1)d = b a3 = a + (2)d = c a + d = b a + 2d = c -d = b + c d = c -b a1 = a^2 a2 = a^2(r)^1 = b^2 a3 = a^2(r)^2 = c^2 ra^2 = b^2 a^2(r^2) = c^2 b = a sqrt (r) c = ar d = ar - a sqrt (r) d = a(r - sqrt r) a = d/(r- sqrt r) c - 2d = d/(r - sqrt r) ar - 2d = d/(r - sqrt r) okay...i give up now...i'm going in circles @_@ :P

OpenStudy (anonymous):

400 :P

OpenStudy (anonymous):

Not right experimentX

OpenStudy (anonymous):

You didnt say i was wrong :D

OpenStudy (callisto):

see right hand side

OpenStudy (anonymous):

@Ishaan94 I was thinking about you when I posted this, try it out man :)

OpenStudy (lgbasallote):

is callisto right?

OpenStudy (anonymous):

callisto's approach is right, however I would appreciate the surd as answer ;)

OpenStudy (experimentx):

1/2*(1-sqrt(2))

OpenStudy (anonymous):

Right ExperimentX

OpenStudy (lgbasallote):

oh...i was soooo close to that! lol =)))

OpenStudy (experimentx):

hahah a+b+c=3/2 2b = a+c and I never considered the probability that b^2 = -ac (damn ...!)

OpenStudy (callisto):

sorry.... i can't help using the calculator... I'll leave it in the surd form next time

OpenStudy (anonymous):

I got the same but I used calculator for calculations :/